OpenStudy (anonymous):

Write the equation of the line with a slope of -1/3 and passing through the point (6, -4).

OpenStudy (camerondoherty):

Ummm y = 1/3x - 4 I think... Not 100% sure though Its in SLope Intercept Form btw Y=Mx+B

OpenStudy (anonymous):

\[y-y _{1}=m(x-x _{1})\]

OpenStudy (shadowlegendx):

y = mx + b y = -1/3x - 4

OpenStudy (shadowlegendx):

This is the equation of the line as it shows the slope, -1/3 and also the y-intercept, -4. This equations tells you where you need to start on the y axis if you are to graph this line and at what rise/run the line will be at.

OpenStudy (anonymous):

I get it now, and thank you. Can you help me with a couple more that are similar to this one?

OpenStudy (anonymous):

\[y-(-4)=\frac{ -1 }{ 3 }\left( x-6) \right)\] or 3(y+4)=-1(x-6) 3y+12=-x+6 3y=-x+6-12 3y=-x-6 \[y=-\frac{ 1 }{ 3 }x-2\]

OpenStudy (anonymous):

put x=6,y=-4 \[-4=\frac{ -1 }{ 3 }*6-2=-2-2~or~-4=-4\] which is true . Hence point (6,-4)lies on this line and slope is -1/3